# Investigating Modality Contribution in Audio LLMs for Music

### v1

@@ 1 Introduction
Similar modality-imbalance issues have been reported in Vision LLMs, where unimodal models can often achieve comparable accuracy to a multimodal one on a vision-language task [12], suggesting unimodal collapse. To better understand this phenomenon without relying on performance metrics, [13, 14] proposed MM-SHAP. They utilize Shapley values [15], a post-hoc explainability technique to quantify the relative contribution of each modality in a model’s prediction, irrespective of its accuracy. We investigate MM-SHAP’s adaptability to the modality of audio. Particularly, our main contributions are that we:
@@ 2 Related Work
Within explainability techniques in machine learning, there is a family of “post-hoc” methods whose aim is to analyze how each input feature contributes to a given model output via input perturbations. These include approaches like LIME (Local Interpretable Model-agnostic Explanations) [16], that explain a prediction by approximating it locally with an interpretable model (e.g., linear regression or a decision tree). To the best of our knowledge, most of the post-hoc techniques applied to music are limited to LIME and LIME-variations [17],[18]. Recently, [19] expanded LIME methods to investigate music understanding models by analyzing the most important features for genre classification from both text and audio modalities. As for modality contribution itself, the DIsentangled Multimodal Explanations (DIME) [20] was proposed as a visualization technique to gain insights into each modality’s contribution to the model’s output. DIME disentangles a model into unimodal and multimodal contributions and then produces visualizations on each sub-model using LIME. Because of its focus on visual explanations, DIME is hard to use as a quantitative measure of modality contribution.
@@
To address the limitations of ablation studies and provide stronger theoretical guarantees than surrogate models like LIME, another family of “post-hoc” methods based on Shapley Values has gained prominence. Shapley Values are the only feature attribution method that satisfies theoretical guarantees [22]. To the best of our knowledge, the only exploration of Shapley values and music audio was performed in [23]. Recently, Parcalabescu and Frank proposed MultiModal-SHAP (MM-SHAP) [13], which quantifies modality contribution in text-vision models by adapting Shapley values to measure how input perturbations across modalities affect the model’s output.
@@ 3.1 Shapley Values and Feature Contribution
where $S$ represents subsets of features excluding the feature $j$, $f_{t}(S)$ is the model’s output for the subset $S$ and token $t$, $n$ defines the total number of features. Therefore, $\phi$ represents how much feature $j$ contributes to the output token $t$. The normalizing factor $\gamma=\frac{n!}{|S|!(n-|S|)!}$ weights the marginal contribution of each feature fairly across all possible subsets of features. A subset’s marginal contribution is weighted proportionally to how often that subset appears when averaging over all permutations. Shapley values satisfy the theorical guarantees of symmetry, additivity, dummy, and additivity, which together constitute a fair payout [22].
The exact computation of a Shapley value requires evaluating all possible subsets of features. As the number of features grows, the computation of Equation 1 has an exponential time complexity of $O(2^{n})$. This means that for most machine learning models, which have large numbers of features, the exact computation of Shapley values is too computationally expensive. There are various approximations of Shapley values proposed to mitigate this, such as Monte-Carlo sampling [25] or KernelSHAP, TreeSHAP, and PermutationSHAP, introduced by [24] in the SHAP framework. In this work, we use PermutationSHAP, which approximates Shapley values by averaging each feature’s marginal contribution across random permutations, measuring the change in the model’s output when the feature is added to preceding subsets. When estimated using $m$ randomly sampled subsets, PermutationSHAP has a complexity of O(mn)O(m*n). We select the default value of $m=10$ from the shap library [24].
@@ 4 Experiments
Models Our experiments compare two Audio LLMs, which represent the best and second-worst performing models on MuChoMusic. MU-LLaMA [3] uses MERT-v1-330M as the audio encoder and LLaMA 2 7B as the language encoder, then trains an adapter to map the MERT output to the language model. All audio is resampled to 24 kHz before encoding. The model is trained with the MusicQA dataset, which is compiled from the MusicCaps [8] and MagnaTagATune [27] datasets. Qwen-Audio [2] uses Qwen-7B to initialize its LLM and Whisper-large-v2 as its base audio encoder. All audio is resampled to 16 kHz prior to encoding. We use Qwen-Audio-Chat for our experiments as it is fine-tuned to answer questions.
@@
For each task and model, we report A-SHAP scores and accuracy, calculated with the code provided in the MuChoMusic GitHub repository. We kept all system instructions with their default values (“You’re a reliable assistant, follow these instructions.”). Unlike [13], we did not mask the instructions, only the question itself.
We also performed a qualitative analysis to investigate how the Shapley values behave for questions requiring single sound event vs. those requiring longer-term audio understanding to generate the answer. We looked at each model and each experiment, filtered only questions with correct answers and selected 5 samples with the highest A-SHAP values, and 5 with the lowest. We removed duplicates, a sample with corrupted audio, and a question with no correct distractor, resulting in 35 unique Q&A. From this set, we first identified questions answered by a single sound event. For these, we manually annotated the relevant audio segment, as in Figure 2. We observed that most of these questions still required a long-term understanding of the audio. To find more single-sound-event examples, we searched for keywords that were common in such questions (e.g., ”interrupted,” ”at the beginning,” ”sound at the end”). We identified a total of 13 Q&A pairs. All of our experiments and code are available on our GitHub [^2] https://github.com/giovana-morais/2025_investigating_mmshap.
@@ 5 Results and Discussion
Table 1 shows that the more accurate model, Qwen-Audio, relies less on audio, whereas MU-LLaMA uses both modalities in a balanced manner. Indicating that there is no relationship between accuracy and modality usage. This was further reinforced by analyzing A-SHAP values grouped by correct and incorrect questions, which showed no trend. Examining questions broken down by knowledge dimension according to MuChoMusic categories (e.g. knowledge, reasoning) also showed no significant effect in the A-SHAP value. We think that this lack of difference is consistent with the findings in [28], where the authors show that LLMs perform as good on MuChoMusic as Audio LLMs, indicating that the audio modality is not leveraged as expected in the answering process. They show that this is caused by MuChoMusic approach to designing the multiple-choice distractors which, as a result, any LLM capable of eliminating the two unrelated answers can achieve an accuracy of 50% by randomly selecting the answer from the two remaining options. What the results of [28] suggest is that what MuChoMusic is, in fact, testing the reasoning skills of an LLM model instead of the perceptual capabilities of Audio LLMs.
@@
As we have seen little change in the audio usage between correct and incorrect examples, we decided to further explore the positive and negative contributions ($\phi_{j,t}$), together with the absolute value that is in fact used to compute MM-SHAP. We analyze the songs as described in Section 4. From the qualitative analysis, we found that in Qwen-Audio the audio features have much lower magnitude than text features, which aligns with the low audio usage. But this does not mean that the audio is being discarded. For each specific token $t\in T$, we see that the positive contributions often align with the audio region that contains the ground truth, and also with the input token that represents that event, as shown in Figure 2 for the token “bell”. We note that the same behavior was found in the other questions with single-sounding events. It is also possible to see other strong activations, and those are not related to the correct answer (in Figure 2, they represent a piano note). We hypothesize that in such cases, the information is useful for the model’s internal mechanisms, but not necessarily are faithful to the data explanation. In examples that require longer-term explanations, the activations are more spread throughout the audio. In MU-LLaMA, the audio and text features had similar magnitude overall, but we noticed that the audio features with high activations were spread throughout the audio in both long-term and single-sounding questions. Although there were positive contributions in the correct regions, they were not as localized as in Qwen-Audio sample. Also, MU-LLaMA activations showed negative contributions as strong as the positive, which affect the final A-SHAP value that takes into account only the magnitude of the feature importance. We provide additional plots comparing all models and experiments in the project’s demo page.
@@
A primary challenge is the lack of ground truth for qualitative analysis. Our analysis was restricted to questions with easily localizable sound events; annotating questions about melody, instrumentation, or structure is harder, yet those question types are the most frequent when evaluating music understanding systems, which makes reliable annotation an open question by itself. Furthermore, the audio segment size, which is tied to the number of text tokens, may be too small to contain perceptually meaningful information, complicating our cross-experiment comparison. For the audio masking, we zeroed the audio segments, but other alternatives include replacing it with white noise or an average between neighbor segments, as in [23] and [29]. Determining an optimal audio window size and the optimal audio mask are open research problems, especially if accounting for temporal and spectral information. This is especially important because the chosen mask must be a “neutral” representation of the data. While MM-SHAP can help us understand how much audio is used in current models, the tool is not enough to hints us about how the audio is used internally in the model, as it is the case with post-hoc explanation methods.
@@ 6 Conclusion
In this work, we adapted MM-SHAP to measure modality contributions in Audio LLMs, aiming to measure and understand how these models leverage audio and text for perception tasks. Through a systematic evaluation of Qwen-Audio and MU-LLaMA, we discovered that the higher-performing model, Qwen-Audio, relied significantly more on its text modality. While text processing appears to drive overall performance, our qualitative analysis confirms that the audio modality is not discarded. Our findings demonstrate that a balanced modality usage does not guarantee superior performance. This insight challenges the assumption that optimal multimodal models must use all modalities equally. Future work should investigate if this is the result of how MuChoMusic defines its distractors or if this is, in fact, how models are using audio information.

### v2

@@ 1 Introduction
Similar modality-imbalance issues have been reported in Vision LLMs, where unimodal models can often achieve comparable accuracy to a multimodal one on a vision-language task [6], suggesting unimodal collapse. To better understand this phenomenon without relying on performance metrics, [19, 20] proposed MultiModal-SHAP (MM-SHAP). They utilize Shapley values [24], a post-hoc explainability technique to quantify the relative contribution of each modality in a model’s prediction, irrespective of its accuracy. We investigate MM-SHAP’s adaptability to the modality of audio. Particularly, our main contributions are that we:
@@ 2 Related Work
Within explainability techniques in machine learning, there is a family of “post-hoc” methods whose aim is to analyze how each input feature contributes to a given model output via input perturbations. These include approaches like LIME (Local Interpretable Model-agnostic Explanations) [22], that explain a prediction by approximating it locally with an interpretable model (e.g., linear regression or a decision tree). To the best of our knowledge, most of the post-hoc techniques applied to music use LIME or LIME-variations [9],[17]. Recently, [25] expanded LIME methods to investigate music understanding models by analyzing the most important features for genre classification from both text and audio modalities. As for modality contribution, the DIsentangled Multimodal Explanations (DIME) [15] was proposed as a visualization technique to gain insights into each modality’s contribution to the model’s output. DIME disentangles a model into unimodal and multimodal contributions and then produces visualizations on each sub-model using LIME. Because of its focus on visual explanations, DIME is hard to use as a quantitative measure of modality contribution.
@@
To address the limitations of ablation studies and provide stronger theoretical guarantees than surrogate models like LIME, another family of “post-hoc” methods based on Shapley values has gained prominence. Shapley values are the only feature attribution method that satisfies theoretical guarantees [18]. To the best of our knowledge, the only exploration of Shapley values and music audio was performed in [2]. Recently, Parcalabescu and Frank proposed MM-SHAP [19], which quantifies modality contribution in text-vision models by adapting Shapley values to measure how input perturbations across modalities affect the model’s output.
@@ 3.1 Shapley Values and Feature Contribution
where $S$ represents subsets of features excluding the feature $j$, $f_{t}(S)$ is the model’s output for the subset $S$ and token $t$, $n$ defines the total number of features. Therefore, $\phi_{j,t}$ represents how much feature $j$ contributes to the output token $t$. The normalizing factor $\gamma=\frac{n!}{|S|!(n-|S|)!}$ weights the marginal contribution of each feature fairly across all possible subsets of features. A subset’s marginal contribution is weighted proportionally to how often that subset appears when averaging over all permutations. Shapley values satisfy the theorical guarantees of symmetry, efficiency, dummy, and additivity, which together constitute a fair payout [18].
The exact computation of Shapley values requires evaluating all possible subsets of features. However, as the number of features grows, the computation of Equation 1 has an exponential time complexity of $O(2^{n})$. This means that, for machine learning models with a large numbers of features, the exact computation of Shapley values is too expensive. In practice, Shapley values are approximated by sampling feature subsets, e.g., using KernelSHAP or PermutationSHAP from the SHAP framework [14]. In this work, we use PermutationSHAP, which approximates Shapley values by averaging each feature’s marginal contribution across random permutations, measuring the change in the model’s output when the feature is added to preceding subsets. When estimated using $m$ randomly sampled subsets, PermutationSHAP has a complexity of O(mn)O(m*n). We use the default value of $m=10$ from the shap library.
@@ 4 Experiments
Models Our experiments compare two Audio LLMs, which represent the best and second-worst performing models on MuChoMusic. MU-LLaMA [13] uses MERT-v1-330M as the audio encoder and LLaMA 2 7B as the language encoder, then trains an adapter to map the MERT output to the language model. All audio is resampled to 24 kHz before encoding. The model is trained with the MusicQA dataset, which is compiled from the MusicCaps [1] and MagnaTagATune [11] datasets. Qwen-Audio [3] uses Whisper-large-v2 as its base audio encoder and Qwen-7B as its base LLM. Audio is resampled to 16 kHz prior to encoding. We use Qwen-Audio-Chat for our experiments as it is fine-tuned to answer questions.
@@
For each task and model, we report A-SHAP scores and accuracy, calculated with the code provided in the MuChoMusic GitHub repository. We kept all system instructions with their default values (“You’re a reliable assistant, follow these instructions.”). Unlike [19], we did not mask system instructions, only the question itself.
We also performed a qualitative analysis to investigate how the Shapley values behave for questions requiring single sound event vs. those requiring longer-term audio understanding to generate the answer. We looked at each model and each experiment, filtered only questions with correct answers and selected 5 samples with the highest A-SHAP values, and 5 with the lowest. We removed duplicates, a sample with corrupted audio, and a question with no correct distractor, resulting in 35 unique Q&A. From this set, we first identified questions answered by a single sound event. For these, we manually annotated the relevant audio segment, as in Figure 2. We observed that most of these questions still required a long-term understanding of the audio. To find more single-sound-event examples, we searched for keywords that were common in such questions (e.g., “interrupted”, ”at the beginning“, ”sound at the end“). We identified a total of 13 Q&A pairs. All of our experiments and code are available on our GitHub [^2] https://github.com/giovana-morais/2025_investigating_mmshap.
@@ 5 Results and Discussion
Table 1 shows that the more accurate model, Qwen-Audio, relies less on audio, whereas MU-LLaMA uses both modalities in a balanced manner. Indicating that there is no relationship between accuracy and modality usage. This was further reinforced by analyzing A-SHAP values grouped by correct and incorrect questions, which showed no trend. Examining questions broken down by knowledge dimension according to MuChoMusic categories (e.g. knowledge, reasoning) also showed no significant effect in the A-SHAP value. We think that this lack of difference is consistent with the findings in [28], where the authors show that LLMs perform as good on MuChoMusic as Audio LLMs, indicating that the audio modality is not leveraged as expected in the answering process. They show that this is caused by MuChoMusic approach to designing the multiple-choice distractors which, as a result, any LLM capable of eliminating the two unrelated answers can achieve an accuracy of 50% by randomly selecting the answer from the two remaining options. What the results of [28] suggest is that what MuChoMusic is, in fact, testing the reasoning skills of an LLM model instead of the perceptual capabilities of Audio LLMs. To verify whether the low audio usage is extended to tasks other than multiple-choice, we additionally evaluated the best-performing model (Qwen-Audio) by asking it to describe the audio clips, with the same number of audio tokens as MC-PI, for fairness.The average A-SHAP increased to 0.73%, suggesting stronger audio usage and significant task influence.
@@
As we have seen little change in the audio usage between correct and incorrect examples, we decided to further explore the positive and negative contributions ($\phi_{j,t}$), together with the absolute value that is in fact used to compute MM-SHAP. We analyze the songs as described in Section 4. From the qualitative analysis, we found that in Qwen-Audio the audio features have much lower magnitude than text features, which aligns with the low audio usage. But this does not mean that the audio is being discarded. For each specific token $t\in T$, we see that the positive contributions often align with the audio region that contains the ground truth, and also with the input token that represents that event, as shown in Figure 2 for the token “bell”. We note that the same behavior was found in the other questions with single-sounding events. It is also possible to see other strong activations, and those are not related to the correct answer (in Figure 2, they represent a piano note). We hypothesize that in such cases, the information is useful for the model’s internal mechanisms, but not necessarily are plausible to the data explanation. In examples that require longer-term explanations, the activations are spread throughout the audio. In MU-LLaMA, the audio and text features had similar magnitude overall, but we noticed that the audio features with high activations were spread throughout the audio in long-term and single-sounding questions. Although there were positive contributions in the correct regions, they were not as concentrated as in Qwen-Audio. Also, MU-LLaMA activations showed negative contributions as strong as the positive, which affect the final A-SHAP value that takes into account only the magnitude of the feature importance. We provide additional interactive plots comparing all models and experiments in the project’s demo page.
@@
A primary challenge is the lack of ground truth for qualitative analysis. Our analysis was restricted to questions with easily localizable sound events; annotating questions about melody, instrumentation, or structure is harder, yet those question types are the most frequent when evaluating music understanding systems, which makes reliable annotation an open question by itself. Furthermore, the audio segment size, which is tied to the number of text tokens, may be too small to contain perceptually meaningful information, complicating our cross-experiment comparison. For the audio masking, we zeroed the audio segments, but other alternatives include replacing them with white noise or an average between neighbor segments, as in [2] and [16]. Determining an optimal audio window size and the optimal audio mask are open research problems, especially if accounting for temporal and spectral information. This is especially important because the chosen mask must be a “neutral” representation of the data. While MM-SHAP and Shapley values can help us understand how much audio is used in current models, the tool is not enough to hint us about how the audio is used internally in the model, as is the case with post-hoc explanation methods.
@@ 6 Conclusion
In this work, we adapt MM-SHAP to measure modality contributions in Audio LLMs, analyzing how audio and text are used in perception tasks. Through a systematic evaluation of Qwen-Audio and MU-LLaMA, we discovered that the best-performing model, Qwen-Audio, relied significantly more on its text modality. While text processing appears to drive overall performance, our qualitative analysis confirms that the audio modality is not discarded, suggesting a gap between listening to the audio and reasoning about it. Our findings also suggest that balanced modality use does not guarantee superior performance, challenging the assumption that optimal multimodal models must use all modalities equally. Future work will systematically study, at scale, how audio information is leveraged, how it survives multimodal integration, and how task formulation (e.g., multiple choice vs. descriptive) influences its use.
7 Acknowledgments
This work is supported by the National Science Foundation under Grant Numbers 2504642 and 2504643. This work was supported in part through the NYU IT High Performance Computing resources, services, and staff expertise.